Use Cramer’s rule to solve for y without solving for x, z and w in the system

2w + x + y + z = 3

−8w − 7x − 3y + 5z = −3

w + 4x + y + z = 6

w + 3x + 7y − z = 1


1
Expert's answer
2021-05-27T05:58:28-0400

2w + x + y + z = 3

−8w − 7x − 3y + 5z = −3

w + 4x + y + z = 6

w + 3x + 7y − z = 1



Let D = "\begin{vmatrix}
 2 & 1 & 1 & 1 \\
 -8 & -7 & -3 & 5\\
 1&4&1&1\\
 1&3&7&-1\\
\end{vmatrix}"


C = "\begin{vmatrix}
 3 \\
 -3 \\
 6 \\
 1 
\end{vmatrix}"


We have to solve for y,


By using Cramer's Rule we have


y = "\dfrac{D_y}{D}"



we can find Dy by replacing the 3rd column of D matrix by the solution Matrix C. after doing this, we get

Dy as


Dy = "\begin{vmatrix}
 2 & 1 & 3 & 1\\
 -8 & -7 & -3 & 5\\
1 & 4 & 6 & 1\\
1 & 3 & 1& -1\\
\end{vmatrix}"



On solving, we get D = 424 and Dy = -78


hence y = -78/424


y = -0.184